Suppose 20 chips numbered 1-20 are put in abox.. Suppose a chip is drawn with replacement repeatedly. What is the probability that a prime (2,3,5,7) will be first taken on the 7th draw?
Answer by stanbon(75887)

Suppose 20 chips numbered 1-20 are put in a box.. Suppose a chip is drawn WITH REPLACEMENT repeatedly.
What is the probability that a prime (2,3,5,7.11,13,17,19) will be first taken on the 7th draw?
------
Pattern::: nnnnnnp
There are 8 primes and 12 non-prime numbers
---
P(1st prime on 7th draw) = (12/20)^6*(8/20) = (3/5)^6*(2/5) = 0.0187
================
Cheers,
Stan H.
-------------------
-----